Calculate the maximum mass of NaIO3 that can be dissolved in 3.00L of 0.200M AgNO3 solution without forming a precipitate
1
Expert's answer
2019-10-28T07:49:47-0400

Answer on Question #97473 – Chemistry – Other


Task:

Calculate the maximum mass of NaIO3 that can be dissolved in 3.00L of 0.200M AgNO3 solution without forming a precipitate.


Solution:

NaIO3 + AgNO3 = AgIO3 + NaNO3

1) n(NaIO3) = n(AgNO3);

n(NaIO3) = C(AgNO3)*V(AgNO3) = 0.200M * 3.00L = 0.600 mol;

2) n(NaIO3) = m(NaIO3) / M(NaIO3);

M(NaIO3) = Ar(Na) + Ar(I) + 3*Ar(O) = 22.9898 + 126.9045 + 3*15.999 = 197.8913 (g/mol).

3) m(NaIO3) = n(NaIO3)*M(NaIO3) = 0.600 mol * 197.8913 g/mol = 118.7348 g.

mmax(NaIO3) = 118.7348 g.


Answer: mmax(NaIO3) = 118.7348 g.